OpenStudy (carissa15):

Have an assignment due tonight and stuck on some questions. Please help? I am not sure of how to balance an equation that has a negative exponent?

OpenStudy (biohazard9064):

I believe that you simplify it.

OpenStudy (carissa15):

I have taken the second derivative, and now trying to find the zeroes of \[2+2x^-3\]

OpenStudy (carissa15):

so obviously will then become\[2x^-3=-2\]

OpenStudy (carissa15):

if I simplify the equation, then I get\[\frac{ 2 }{ x^3 }+2\]

OpenStudy (carissa15):

but still not sure what to do to "get rid" of the exponent to leave x on it's own and find the value of x?

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

what is your question?

OpenStudy (welshfella):

2 x^-3 = -2 divide both sides by 2 x^-3 = -1 1 / x^3 = -1 cross multiply - x^3 = 1 divide both sides by -1 x^3 = -1 x = -1

OpenStudy (carissa15):

great thank you